OpenStudy (anonymous):

Equilateral triangle RST is similar to the equilateral triangle DEF with the ratio of similarity 26. Which is equivalent to 3RS + 5ST over 6DE + 7EF 16 14 12 8

OpenStudy (anonymous):

is that the full question?

OpenStudy (anonymous):

yea

myininaya (myininaya):

what is the ratio?

OpenStudy (anonymous):

i gues there is smthing missing

OpenStudy (anonymous):

It's equilateral, so RS=ST (call this a) and DE=EF (call this b) a=26b (3a+5a)/(6b+7b) =8a/(13b) =8(26b)/13b =8*2*13b/(13b) =16